OpenStudy (seattle12345):

Please help me to solve this: You have one type of nut that sells for $1.90/lb and another type of nut that sells for $6.60/lb. You would like to have 28.2 lbs of a nut mixture that sells for $5.60/lb. How much of each nut will you need to obtain the desired mixture?

OpenStudy (pfenn1):

Let n_1 be one type of nut and n_2 be the other. \[1.90n_1+6.60n_2=(5.60)(28.2)\]\[n_1+n_2=28.2\]\[n_2=28.2-n_1\]\[1.90n_1+6.60(28.2-n_1)=(5.60)(28.2)\]Rearrange and solve for n_1.

OpenStudy (seattle12345):

I did that and came up with -4.7\[n _{1}\]=28.92

OpenStudy (pfenn1):

That isn't the answer I got.

OpenStudy (seattle12345):

whoops did it again and came up with -4.7n=-28.2... which equals 6

OpenStudy (pfenn1):

Yes, that is what I got.

OpenStudy (seattle12345):

then \[n_{2}=22.2\]?

OpenStudy (pfenn1):

Yes.